OpenStudy (anonymous):

If 3/4 = x/8, then x is equal to ?

OpenStudy (coolguy34):

6 is your answer.

OpenStudy (isaidavila):

8*3=4*x

OpenStudy (anonymous):

Basically 3/4 is = ?/8 To get 8 from 3/4 we have to multiply 4 by 2.

OpenStudy (anonymous):